OpenStudy (anonymous):

If the endpoints of the diameter of a circle are (-10, 12) and (12, 18), what are the coordinates of the center of the circle?

OpenStudy (agent0smith):

The centre will be halfway between those two points, since the centre is halfway along the diameter. Use the midpoint formula to find it \[\Large \left( \frac{ x_1+x_2 }{ 2 }, \frac{ y_1+y_2 }{ 2 } \right)\]

OpenStudy (anonymous):

*facepalm* thank you